the inverse demand functions of two towns for a product are given as ( Q1/2)1/4=P and {2-Q2/10}=P}= P find the market demand of these two towns
1
Expert's answer
2020-03-27T10:19:43-0400

a) (Q1/2)^1/4 = P or Qd1 = 2P^4,

2 - Q2/10 = P or Qd2 = 20 - 10P,

The market demand of these two towns is the sum of individual demands, so:

"Qd = Qd1 + Qd2"

"Qd = 2P^4 - 10P + 20."